OpenStudy (konradzuse):

Find the area of the region. Inside r = 2a cos(θ) and outside r = a

OpenStudy (anonymous):

\[\int_{0}^{2\pi}\int_{2a * cos(\theta)}^{a} r dr d\theta\]

OpenStudy (konradzuse):

????

OpenStudy (anonymous):

what??

OpenStudy (konradzuse):

What does that even mean?

OpenStudy (anonymous):

that's double integral

OpenStudy (konradzuse):

how do you go about that? Never have seen that form before....

OpenStudy (konradzuse):

@experimentX

OpenStudy (anonymous):

\[\int_{0}^{2\pi}\int_{2a * cos(\theta)}^{a} r dr d\theta\] \[\int_{0}^{2\pi}|_{2a * cos(\theta)}^{a} r^2/2 d\theta\] \[\int_{0}^{2\pi}(a)^2-(2a cos(\theta))^2/2 d\theta\]

OpenStudy (anonymous):

you integrate it again, this time w.r.t theta

OpenStudy (konradzuse):

hmm

OpenStudy (konradzuse):

wow OS crashed hard...

OpenStudy (anonymous):